OpenStudy (anonymous):

\[\sum_{-6}^{6} (k^2+2k) =?\]

OpenStudy (anonymous):

$$\sum_{k=-6}^6\left(k^2+k\right)=\sum_{k=-6}^6k^2+\sum_{k=-6}^6k=2\sum_{k=1}^6k^2=\dots$$

OpenStudy (anonymous):

@oldrin.bataku Thank you so much

OpenStudy (anonymous):

Do you see what we did? \(\sum\limits_{k=-n}^n k=0\) is trivial, and we know that \(\sum\limits_{k=-n}^0k^2=\sum\limits_{k=0}^nk^2\) which made it simple.

OpenStudy (anonymous):

Abselutely get it. I appreciate your effort . Thanks